a hiking club charges an initial fee of $60 and then $12 for each hinking trip. write an expression to model the total cost and then evaluate the expression for 1 to 6 hikes.
Answer by askmemath(368)

You pay $60 as membership fee
$12 per hike
Let's say you went for 3 hikes in a year
Then your total cost = 60+12x3 = $96
4 hikes
60+12x4 = $108
So your expression bcomes 60+12X where X is the number of hikes you take.